OpenStudy (anonymous):

how do I do I state the order of the following maths differential equation. (d to the power 3 times y divide by d times x to the power 3) to the square root 2.

OpenStudy (zehanz):

It is \(\left(\dfrac{d^3y}{dx^3}\right)^{\sqrt{2}}\). The \(\sqrt{2}\) is just a calculation that has to be done with \(\dfrac{d^3y}{dx^3}\), so the order must be 3.